OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 4x + 5y = -9.

OpenStudy (imstuck):

get this into y = mx + b form. Solve it for y, and the value that is with the x after all is said and done is the slope.

OpenStudy (anonymous):

set the equation to y = mx+b where m is the slope and b is the y-int.

OpenStudy (imstuck):

The y intercept will be the number in the "b" position in your y=mx+b equation.

OpenStudy (imstuck):

Can you solve for y?

OpenStudy (anonymous):

I deleted it

OpenStudy (imstuck):

Thank you for deleting that.